OpenStudy (anonymous):

Having a bit of trouble with \[\int_0^{\pi/2}\ln(\sin(x))dx\]. I'll post the solution given and the part that confuses me. Any help would be much appreciated!

OpenStudy (anonymous):

Here's the given solution.

OpenStudy (anonymous):

What I don't get is why \[\int_0^{\pi/2}\ln(\sin(x))dx = \int_0^{\pi/2}\ln(\cos(x))dx\]when it's clearly stated that this was achieved using \(x = \frac{\pi}{2} - u\). \[\int_0^{\pi/2}\ln(\cos(x))dx = \int_0^{\pi/2}\ln(\cos(u))du\]just doesn't seem to work, as I've seen on other proofs of this on other sites.

OpenStudy (anonymous):

NOTE: That's what it meant by recommended substitution on the actual problem.

OpenStudy (anonymous):

oooh i didn't get it at all!

OpenStudy (anonymous):

now i see very clever

OpenStudy (anonymous):

put \(x=\frac{\pi}{2}-u\) \(du=-dx\) and also \(u(0)=\frac{\pi}{2}, u(\frac{\pi}{2})=0\) you get that the integrals are the same

OpenStudy (anonymous):

x=pi/2-u dx=-du integrate(ln(sin(x))dx) from x=0 to x=pi/2 =integrate(ln(cos(u))*(-du)) from x=0 to x=pi/2 =-integrate(ln(cos(u)) from u=pi/2 to u=0 =integrate(ln(cos(u)) from u=0 to u=pi/2

OpenStudy (anonymous):

Oh! I think that I get it now. Is because the boundaries are specifically 0 and pi/2?

OpenStudy (anonymous):

right, otherwise it would not work

OpenStudy (anonymous):

Ah ok. Thank you so much! Just been bothering me for like an hour now. :)

OpenStudy (anonymous):

yeah, without that hint it bothered me too